OpenStudy (anonymous):

please help! how do you solve log base2 (x+1) - log base2 (x-4)=3

Parth (parthkohli):

One thing: \[\log(a) - \log(b) = \log(a/b)\]

Parth (parthkohli):

Also, log is not defined for a negative or zero. So keep that in mind.

OpenStudy (anonymous):

\[\log_{2} \frac{ x+1 }{ x-4 }=3 \]

OpenStudy (anonymous):

then do i put 2 to the third power?

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

i got X= 4.7142 is that correct

Parth (parthkohli):

Yeah.

OpenStudy (anonymous):

ok thank u!!!!